How do I solve for a in the problem 4a+2=3b-5 ?
Found 2 solutions by addingup, Edwin McCravy:
Answer by addingup(3677)

4a = 3b-5-2
4a = 3b-7
a = (3b-7)4

Answer by Edwin McCravy(20056)


That 4 in that last step in the solution above
should be divided, not multiplied.

Here's the solution in more detail: 


4a+2 = 3b-5

Get rid of the +2 on the left by adding -2 to both sides:

4a+2 = 3b-5
  -2     -2
-----------
4a+0 = 3b-7

Ignore the 0:

4a   = 3b-7

Get rid of the multiplied 4 by dividing both sides by 4:

 4a    3b-7
覧覧 = 覧覧
  4     4

Cancel the 4's on the left

 4a    3b-7
覧覧 = 覧覧
  4     4
 
       3b-7
  a  = 覧覧
        4